# Stony Brook, NY 11794, USA

### v1

@@ 1 Introduction and summary
Under the $U(1)_{r}$-twisted reduction with $\gamma\geq 1$, some Coulomb branches are lifted, and the UV 4d AD theories flow to IR 3d theories with simpler Coulomb branches. When the 4d AD theory has no Higgs branch and the $U(1)_{r}$-twisted reduction lifts all the Coulomb branches, we expect that the resulting 3d theories flow to either rank-0 SCFTs or (unitary) topological field theories (TQFTs), both of which have trivial Coulomb and Higgs branches. This scenario occurs for the $U(1)_{r}$-twisted reduction of $(A_{1},A_{2n})$ AD theories with $\gamma$ coprime to the $N$, which is $(2n+3)$ in the case. In the TQFT cases, extended supersymmetry is absent, as it acts trivially on all IR observables, which are non-local operators. The TQFT supports a unitary rational VOA at the boundary. In the the rank-0 SCFT cases, A-twisting results in a (semisimple) non-unitary TQFT that supports a non-unitary rational VOA at the boundary. Interestingly, unitary and non-unitary TQFTs obtained through the $U(1)_{r}$-twisted reduction (and subsequent topological twisting) of a 4d AD theory for different values of $\gamma$ are interconnected by a Galois conjugation operation Dedushenko:2018bpp ; ArabiArdehali:2024ysy . Thus, the $U(1)_{r}$-twisted reductions provide a physical realization of the Galois conjugation.
@@
The reduction procedure has been streamlined in ArabiArdehali:2024ysy for theories with a Lagrangian description, involving a reductive gauge group $G$ of rank $r_{G}$ and manifesting 4d $\mathcal{N}=1$ supersymmetry.
@@
  • In the first step, one looks for stationary loci of a piecewise quadratic function $Q^{\gamma}\in C^{1}(\mathfrak{h}_{\text{cl}})$ (determined by the gauge and $U(1)_{r}$ charges of the matter fields—see (17)) on the moduli-space $\mathfrak{h}_{\text{cl}}=(-\frac{1}{2},\frac{1}{2}]^{r_{G}}$ of the gauge holonomies around the circle. If these loci are isolated, they correspond to the 3d vacua. If the stationary loci are extended, however, they signal gauge invariant BPS monopoles in the 3d theory. These monopole operators parametrize the perturbatively unlifted 3d Coulomb branch.
@@
  • In the second step, one examines the slopes of various piecewise linear functions $L_{F}^{\gamma}\in C^{0}(\mathfrak{h}_{\text{cl}})$ (determined by the global $F$ as well as the gauge and $U(1)_{r}$ charges of the various fields—see e.g. (18), (25)), along the flat directions of $Q^{\gamma}$. These slopes give the global (possibly R-) symmetry charges of the corresponding BPS monopoles. If one finds monopoles of R-charge $2$ and all other global charges $0,$ they ought to be included in the superpotential of the 3d theory as non-perturbatively generated corrections. [^3] That monopole superpotentials can be generated in the circle reduction of 4d $\mathcal{N}=1$ theories to 3d $\mathcal{N}=2$ has been discussed in detail in Aharony:2013dha ; Aharony:2013kma , as well as ArabiArdehali:2019zac where a holonomy saddle scenario similar to the ones in this work was treated. One way to understand this is that the monopole superpotentials break in 3d certain abelian symmetries that were anomalous in 4d (equivalently, the condition of exact marginality of the monopole in 3d is equivalent to the one of the anomaly cancellation in 4d). See also Nardoni:2024sos for a recent discussion on the fate of these anomalies in the compactifications of 4d theories to 3d and 2d. These would then further shrink the 3d Coulomb branch, from the stationary loci of $Q^{\gamma}$ to a smaller subset. In all the examples treated in ArabiArdehali:2024ysy and the present paper, these smaller subsets are isolated. Therefore the 3d EFTs do not have Coulomb branches of vacua.
@@
  • In the third and final step, the massless field content supported on the Coulomb branch of vacua obtained as above, and the various (mixed) Chern-Simons couplings of the 3d EFT, are determined using again the $Q^{\gamma}$ and $L_{F}^{\gamma}$ functions, as reviewed in the next section. In fact, the $U(1)_{r}$-twist would generically imply that all KK modes of the matter fields are massive and integrated out, however this can be compensated by going on a non-generic point of the 3d Coulomb branch where some fields remain massless. The price to pay is that the 4d gauge symmetry is broken to a subgroup in 3d, which is typically the Cartan (however see Appendix A for a non-abelian example).
@@
The resulting 3d EFT has only manifest 3d $\mathcal{N}=2$ supersymmetry, but it either undergoes non-trivial IR supersymmetry enhancement and flow to an $\mathcal{N}=4$ SCFT, or dynamically generates a mass gap and flows to a unitary TQFT. Thus, through the twisted reduction of 4d $\mathcal{N}=2$ SCFTs and their $\mathcal{N}=1$ Lagrangians, one can systematically generate 3d $\mathcal{N}=2$ gauge theories that exhibit either non-trivial IR SUSY enhancement or a dynamically generated mass gap.
@@
In this paper, we consider the 3d $\mathcal{N}=2$ SUSY enhancing gauge theories that arise from the $U(1)_{r}$-twisted circle reduction of the 4d $\mathcal{N}=1$ Lagrangians for the 4d $\mathcal{N}=2$ SCFTs of type $(A_{1},A_{n})$ and $(A_{1},D_{n})$, focusing on $\gamma=1$. The 3d theories resulting from $\gamma=-1$ are obtained via simple conjugations of those arising from $\gamma=1$, as explained in Subsection 3.5. An example with $\gamma=2$ is treated in Appendix A.
From the $(A_{1},A_{2n})$ family, we obtain the Gang-Kim-Stubb $\mathcal{T}_{n}$ theories Gang:2023rei , generalizing the derivation made in the special cases $n=1,2$ in ArabiArdehali:2024ysy . From the remaining three families we obtain abelian SUSY enhancing gauge theories somewhat similar to the Gang-Kim-Stubb family, as follows.
@@
$\mathcal{W}_{V}=V_{1,-1,0,\dots,0}+V_{0,1,-1,0,\dots,0}+\dots+V_{0,\dots,0,1}\qquad\qquad(\text{$n-1$ terms}),$ (3)
where $V_{m_{1},\dots,m_{r_{G}}}$ denotes the BPS monopole with magnetic gauge fluxes $m_{1},\dots,m_{r_{G}}$.
@@
We begin in Section 2 by reviewing the $U(1)_{r}$-twisted 4$\to$3d reduction, given a 4d $\mathcal{N}=1$ gauge theory description of the 4d $\mathcal{N}=2$ SCFT.
In Section 3 we apply the reduction procedure with $\gamma=\pm 1$ to the $(A_{1},A_{n})$ and $(A_{1},D_{n})$ Argyres-Douglas theories. From the $(A_{1},A_{2n})$ family we recover the Gang-Kim-Stubbs theories. From the remaining three families $(A_{1},A_{2n-1})$, $(A_{1},D_{2n+1})$ and $(A_{1},D_{2n})$, three new infinite families of SUSY enhancing 3d gauge theories with monopole superpotentials are derived. The first two members of each of the new families are treated in detail: in most cases their extra SUSY current operators are determined, and in all cases their Higgs and Coulomb branch Hilbert series are matched with the expectations of the reduction from 4d (namely, trivial 3d Coulomb branch and the 3d Higgs branch identical to the 4d one), using appropriate limits of the 3d superconformal index Razamat:2014pta . The simplest new theory is the one arising from $(A_{1},A_{3})$, which we discuss in detail in Subsection 3.2.1. For this we present a dual description with manifest $\mathcal{N}=3$ supersymmetry, obtained by gauging the $SU(2)$ symmetry on the Coulomb branch of the $T[SU(2)]$ theory Intriligator:1996ex ; Gaiotto:2008ak with Chern-Simons level $k=3$ along with a topological sector.
In Section 4, the TQFT arising from A-twisting the simplest new SUSY enhancing theory, namely the one arising from twisted reduction of $(A_{1},A_{3})$, is studied. In particular, its modular data are determined using Bethe roots and BPS surgery techniques and matched with those expected for the corresponding $\widehat{\mathfrak{su}}(2)_{-4/3}$ boundary VOA.
Section 5 outlines some future directions. It also discusses an observation on the possibility of systematically obtaining unitary TQFTs from $\mathcal{N}=2$ preserving superpotential deformations of 3d $\mathcal{N}=4$ SCFTs with trivial Coulomb branch but non-trivial Higgs branch, such as those devised in the present work.
@@
  • Gauge symmetry and field content. Reducing the theory with the $U(1)_{r}$-twist would naively integrate out all the matter fields, since the mass of the KK modes is shifted such that none of them stays massless. However, this is true at generic points of the 3d Coulomb branch, while at special points some fields might becomes massless if the gauge holonomy $\boldsymbol{x}$ compensates the mass induced by the twist. This is indeed the case at the saddle point $\boldsymbol{x}^{\ast}$ we focus on. The price to pay is that, except for specific points like the origin, the gauge group is partially broken on the Coulomb branch. The breaking can be at most to the Cartan subgrop, which will be the preserved gauge symmetry in most of the examples of this paper (except for Appendix A). In order to determine the massless field content of the EFT we simply have to look at which arguments of the $\vartheta$ functions in $L_{R_{s}}^{\gamma}(\boldsymbol{x}^{*})$ vanish. The light vector fields will tell us what is the preserved gauge symmetry, while the light chiral fields will tell us what is the matter content.
  • Chern-Simons couplings. The quadratic function $Q^{\gamma}(\boldsymbol{x})$ encodes the gauge-gauge CS interactions of the low energy theory at the point $\boldsymbol{x}$ of the Coulomb branch. The bare CS levels of the EFT of our interest are then given by the Hessian at the saddle point $K_{ij}=\partial_{i}\partial_{j}Q^{\gamma}(\boldsymbol{x}^{*})$.
@@
Once the 3d $\mathcal{N}=2$ EFT data is determined as above, we can compute monopole gauge and $R_{s}$ charges of the monopole operators via [^4] A few comments are in order regarding these equations. First, we point out that we use an orientation-reversed convention here compared to ArabiArdehali:2024ysy . Second, these equations apply to abelian symmetries; for non-abelian ones they can be applied to each factor of their Cartan subgroups. Finally, similar equations apply also if the symmetry is a global symmetry—see e.g. (26). (see e.g. Borokhov:2002ib ; Borokhov:2002cg ; Borokhov:2003yu ; Gaiotto:2008ak ; Benna:2009xd ; Bashkirov:2010kz ; Cremonesi:2013lqa ; Closset:2016arn ; Pasquetti:2019uop )
@@ 1 $\frac{3}{4n+6}+\frac{4n+3}{4n+6}s$ $e_{n}^{(n)}$ $e_{n}^{(n)}$
Remark. Since we have one fewer monopole superpotential terms (namely $n-1$) than the rank (namely $n$), the superpotential (38) under-determines the vector $k_{jR_{s}}$ of mixed gauge-$R_{s}$ CS couplings. Table 1 needs therefore to be accompanied with at least one entry of $k_{jR}$ as a function of $s$, to provide a complete description. For the case of $(A_{1},A_{2})$, i.e. the Gang-Yamazaki theory Gang:2018huc , the additional required data is found from averaging $\partial_{x}L_{R_{s}}^{\gamma=1}(x)$ on the two sides of the holonomy saddle as in ArabiArdehali:2024ysy to be $k_{gR_{s}}=\frac{s-1}{20}$. For $(A_{1},A_{2n})$ with $n>1$, i.e. the Gang-Kim-Stubbs $\mathcal{T}_{n>1}$ theories, we have similarly computed $k_{g_{n}R_{s}}=\frac{-1+4n(-2+s)-5s}{12+8n}.$ [^6] As a consistency check, we note that for $n=2$ and $s=\tfrac{1}{3}$ we get $k_{g_{2}R_{1/3}}=-\tfrac{4}{7},$ matching the result reported below Eq. (4.2) in ArabiArdehali:2024ysy (modulo the different sign convention, see footnote 4).
@@ 3.2.1 $(A_{1},A_{3})$ in more detail
We propose following IR duality:
@@
Here, the $T[SU(2)]$ theory represents the 3d $\mathcal{N}=4$ SQED with $N_{f}=2$ flavors Intriligator:1996ex ; Gaiotto:2008ak . This theory has both Coulomb and Higgs branches, each of which is isomorphic to $\mathbb{C}^{2}/\mathbb{Z}_{2}$, on which the $SU(2)^{(1)}$ and $SU(2)^{(2)}$ flavor symmetries act respectively. The theory on the right-hand side is obtained by gauging the $SU(2)^{(1)}$ symmetry with $\mathcal{N}=3$ Chern-Simons terms with level $k=3$. The $\mathcal{N}=3$ theory is expected to flow to a $\mathcal{N}=4$ SCFT with a SUSY enhancement, due to the nilpotent property of the moment maps of the $SU(2)^{(1)}$ Gaiotto:2008ak . The enhancement can be explicitly confirmed using the superconformal index Evtikhiev:2017heo . Gauging lifts all the Coulomb branch, leaving the theory with only the Higgs branch $\mathbb{C}^{2}/\mathbb{Z}_{2}$ and $SU(2)$ flavor symmetry, which matches that of the $(A_{1},A_{3})$ EFT. This duality can also be verified through computations of the superconformal index, in particular one can detect the presence of the topological sector by turning on background magnetic fluxes for the $SU(2)$ flavor symmetry Beratto:2021xmn . The round 3-sphere partition functions $Z_{S^{3}}$ for the theories appearing in the duality can be computed using localization, yielding

### v2

@@ 1 Introduction and summary
Under the $U(1)_{r}$-twisted reduction with $\gamma\geq 1$, some Coulomb branches are typically lifted, and the UV 4d AD theories flow to IR 3d theories with simpler Coulomb branches. When the 4d AD theory has no Higgs branch and the $U(1)_{r}$-twisted reduction lifts all the Coulomb branches, we expect that the resulting 3d theories flow to either rank-0 SCFTs or (unitary) topological field theories (TQFTs), both of which have trivial Coulomb and Higgs branches. This scenario occurs for the $U(1)_{r}$-twisted reduction of $(A_{1},A_{2n})$ AD theories with $\gamma$ coprime to the $N$, which is $(2n+3)$ in this case. In the TQFT cases, extended supersymmetry is absent, as it acts trivially on all IR observables, which are non-local operators. The TQFT supports a unitary rational VOA at the boundary. In the the rank-0 SCFT cases, A-twisting results in a (semisimple) non-unitary TQFT that supports a non-unitary rational VOA at the boundary. Interestingly, unitary and non-unitary TQFTs obtained through the $U(1)_{r}$-twisted reduction (and subsequent topological twisting) of a 4d AD theory for different values of $\gamma$ are interconnected by a Galois conjugation operation Dedushenko:2018bpp ; ArabiArdehali:2024ysy . Thus, the $U(1)_{r}$-twisted reductions provide a physical realization of the Galois conjugation.
@@
The reduction procedure has been streamlined in ArabiArdehali:2024ysy for theories with a Lagrangian description, involving a reductive gauge group $G$ of rank $r_{G}$, and manifesting 4d $\mathcal{N}=1$ supersymmetry.
@@
  • In the first step, one looks for stationary loci of a piecewise quadratic function $Q^{\gamma}\in C^{1}(\mathfrak{h}_{\text{cl}})$ (determined by the gauge and $U(1)_{r}$ charges of the matter fields—see (17)) on the moduli-space $\mathfrak{h}_{\text{cl}}=(-\frac{1}{2},\frac{1}{2}]^{r_{G}}$ of the gauge holonomies around the circle. If these loci are isolated, they correspond to the 3d vacua. If the stationary loci are extended, however, they signal gauge invariant BPS monopoles in the 3d theory. These monopole operators parametrize the perturbatively unlifted 3d $\mathcal{N}=2$ Coulomb branch.
@@
  • In the second step, one examines the slopes of various piecewise linear functions $L_{F}^{\gamma}\in C^{0}(\mathfrak{h}_{\text{cl}})$ (determined by the global $F$ as well as the gauge and $U(1)_{r}$ charges of the various fields—see e.g. (18), (25)), along the flat directions of $Q^{\gamma}$. These slopes give the global (possibly R-) symmetry charges of the corresponding BPS monopoles. If one finds monopoles of R-charge $2$ and all other global charges $0,$ they ought to be included in the superpotential of the 3d theory as non-perturbatively generated corrections. [^3] That monopole superpotentials can be generated in the circle reduction of 4d $\mathcal{N}=1$ theories to 3d $\mathcal{N}=2$ has been discussed in detail in Aharony:2013dha ; Aharony:2013kma , as well as ArabiArdehali:2019zac where a holonomy saddle scenario similar to the ones in the present work was treated. One way to understand this is that the monopole superpotentials break in 3d certain abelian symmetries that were anomalous in 4d (equivalently, the condition of exact marginality of the monopole in 3d is equivalent to the one of the anomaly cancellation in 4d). See also Nardoni:2024sos for a recent discussion on the fate of these anomalies in the compactifications of 4d theories to 3d and 2d. These would then further shrink the 3d $\mathcal{N}=2$ Coulomb branch, from the stationary loci of $Q^{\gamma}$ to a smaller subset. In all the examples treated in ArabiArdehali:2024ysy and the present paper, these smaller subsets are isolated. Therefore the 3d EFTs do not have Coulomb branches of vacua.
@@
  • In the third and final step, the massless field content supported on the Coulomb branch of vacua obtained as above, and the various (mixed) Chern-Simons couplings of the 3d EFT, are determined using again the $Q^{\gamma}$ and $L_{F}^{\gamma}$ functions. This is reviewed in the next section. In fact, the $U(1)_{r}$-twist would generically imply that all KK modes of the matter fields are massive and integrated out, however this can be compensated by going on a non-generic point of the classical 3d $\mathcal{N}=2$ Coulomb branch where some fields remain massless. The price to pay is that the 4d gauge symmetry is broken to a subgroup in 3d, which is typically the Cartan (however see Appendix A for a non-abelian example).
@@
The resulting 3d EFT has only manifest 3d $\mathcal{N}=2$ supersymmetry, but it either undergoes non-trivial IR supersymmetry enhancement and flow to an $\mathcal{N}=4$ SCFT, or dynamically generates a mass gap and flows to a unitary TQFT. Thus, through the twisted reduction of 4d $\mathcal{N}=2$ SCFTs using their $\mathcal{N}=1$ Lagrangians, one can systematically generate 3d $\mathcal{N}=2$ gauge theories that exhibit either non-trivial IR SUSY enhancement or a dynamically generated mass gap.
@@
In this paper, we consider the 3d $\mathcal{N}=2$ SUSY enhancing gauge theories that arise from the $U(1)_{r}$-twisted circle reduction of the 4d $\mathcal{N}=1$ Lagrangians for the 4d $\mathcal{N}=2$ SCFTs of type $(A_{1},A_{n})$ and $(A_{1},D_{n})$. We focus on $\gamma=1$. The 3d theories resulting from $\gamma=-1$ are obtained via simple conjugations of those arising from $\gamma=1$, as explained in Subsection 3.5. An example with $\gamma=2$ is treated in Appendix A.
From the $(A_{1},A_{2n})$ family, we obtain the Gang-Kim-Stubb $\mathcal{T}_{n}$ theories Gang:2023rei , generalizing the derivation made in the special cases $n=1,2$ in ArabiArdehali:2024ysy . From the remaining three families, we obtain abelian SUSY enhancing gauge theories somewhat similar to the Gang-Kim-Stubb family, as follows.
@@
Here $V_{m_{1},\dots,m_{r_{G}}}$ denotes the BPS monopole with magnetic gauge fluxes $m_{1},\dots,m_{r_{G}}$.
@@
[…]
In Section 3 we apply the reduction procedure with $\gamma=\pm 1$ to the $(A_{1},A_{n})$ and $(A_{1},D_{n})$ Argyres-Douglas theories. From the $(A_{1},A_{2n})$ family we recover the Gang-Kim-Stubbs theories. From the remaining three families $(A_{1},A_{2n-1})$, $(A_{1},D_{2n+1})$ and $(A_{1},D_{2n})$, three new infinite families of SUSY enhancing 3d gauge theories with monopole superpotentials are derived. The first two members of each new family are treated in detail: in most cases their extra SUSY current operators are determined, and in all cases their Higgs and Coulomb branch Hilbert series are matched with the expectations from the 4d reduction (namely, trivial 3d Coulomb branch and the 3d Higgs branch identical to the 4d one), using appropriate limits of the 3d superconformal index Razamat:2014pta . The simplest new theory is the one arising from $(A_{1},A_{3})$, which we discuss in detail in Subsection 3.2.1. For this we present a dual description with manifest $\mathcal{N}=3$ supersymmetry, obtained by gauging the $SU(2)$ symmetry on the Coulomb branch of the $T[SU(2)]$ theory Intriligator:1996ex ; Gaiotto:2008ak with Chern-Simons level $k=3$, besides a topological sector.
In Section 4, the TQFT arising from A-twisting the simplest new SUSY enhancing theory, namely the one arising from twisted reduction of $(A_{1},A_{3})$, is studied. In particular, its modular data are determined using Bethe roots and BPS surgery techniques, and matched with those expected for the corresponding $\widehat{\mathfrak{su}}(2)_{-4/3}$ boundary VOA.
Section 5 outlines some future directions. It also discusses an observation on the possibility of systematically obtaining unitary TQFTs from $\mathcal{N}=2$ preserving superpotential deformations of 3d $\mathcal{N}=4$ SCFTs with trivial Coulomb branch but non-trivial Higgs branch, such as those constructed in this work.
@@
  • Gauge symmetry and field content. Reducing the theory with the $U(1)_{r}$-twist would naively integrate out all the matter fields, since the mass of the KK modes is shifted such that none of them stays massless. However, this is true at generic points of the 3d Coulomb branch, while at special points some fields might becomes massless if the gauge holonomy $\boldsymbol{x}$ compensates the mass induced by the twist. This is indeed the case at the saddle point $\boldsymbol{x}^{\ast}$ we focus on. The price to pay is that, except for specific points like the origin, the gauge group is partially broken on the Coulomb branch. The breaking can be at most to the Cartan subgroup, which will be the preserved gauge symmetry in most of the examples of this paper (except for Appendix A). In order to determine the massless field content of the EFT we simply have to look at which arguments of the $\vartheta$ functions in $L_{R_{s}}^{\gamma}(\boldsymbol{x}^{*})$ vanish. The light vector fields will tell us what is the preserved gauge symmetry, while the light chiral fields will tell us what is the matter content.
  • Chern-Simons couplings. The quadratic function $Q^{\gamma}(\boldsymbol{x})$ encodes the gauge-gauge CS interactions of the low energy theory at the point $\boldsymbol{x}$ of the Coulomb branch. The bare CS levels of the EFT of our interest are then given by the Hessian at the saddle point [^4] Since we evaluate exactly at $\boldsymbol{x}^{*}$, we need not worry about “inner” versus “outer” patches. See footnote 7 in ArabiArdehali:2024ysy . $K_{ij}=\partial_{i}\partial_{j}Q^{\gamma}(\boldsymbol{x}^{*})$.
@@
Once the 3d $\mathcal{N}=2$ EFT data is determined as above, we can compute the gauge and $R_{s}$ charges of the monopole operators via [^5] A few comments are in order regarding these equations. First, we point out that we use an orientation-reversed convention here compared to ArabiArdehali:2024ysy . Second, these equations apply to abelian symmetries; for non-abelian ones they can be applied to each factor of their Cartan subgroups. Finally, similar equations apply also if the symmetry is a global symmetry—see e.g. (26). (see e.g. Borokhov:2002ib ; Borokhov:2002cg ; Borokhov:2003yu ; Gaiotto:2008ak ; Benna:2009xd ; Bashkirov:2010kz ; Cremonesi:2013lqa ; Closset:2016arn ; Pasquetti:2019uop )
@@ 1 $\frac{3}{4n+6}+\frac{4n+3}{4n+6}s$ $e_{n}^{(n)}$ $e_{n}^{(n)}$
Remark. Since we have one fewer monopole superpotential terms (namely $n-1$) than the rank (namely $n$), the superpotential (38) under-determines the vector $k_{jR_{s}}$ of mixed gauge-$R_{s}$ CS couplings. Table 1 needs therefore to be accompanied with at least one entry of $k_{jR}$ as a function of $s$, to provide a complete description. For the case of $(A_{1},A_{2})$, i.e. the Gang-Yamazaki theory Gang:2018huc , the additional required data is found from averaging $\partial_{x}L_{R_{s}}^{\gamma=1}(x)$ on the two sides of the holonomy saddle as in ArabiArdehali:2024ysy to be $k_{gR_{s}}=\frac{s-1}{20}$. For $(A_{1},A_{2n})$ with $n>1$, i.e. the Gang-Kim-Stubbs $\mathcal{T}_{n>1}$ theories, we have similarly computed $k_{g_{n}R_{s}}=\frac{-1+4n(-2+s)-5s}{12+8n}.$ [^7] As a consistency check, we note that for $n=2$ and $s=\tfrac{1}{3}$ we get $k_{g_{2}R_{1/3}}=-\tfrac{4}{7},$ matching the result reported below Eq. (4.2) in ArabiArdehali:2024ysy (modulo the different sign convention, see footnote 5).
@@ 3.2.1 $(A_{1},A_{3})$ in more detail
We propose the following IR duality:
@@
Here, the $T[SU(2)]$ theory represents the 3d $\mathcal{N}=4$ SQED with $N_{f}=2$ flavors Intriligator:1996ex ; Gaiotto:2008ak . This theory has both Coulomb and Higgs branches, each of which is isomorphic to $\mathbb{C}^{2}/\mathbb{Z}_{2}$, on which the $SU(2)^{(1)}$ and $SU(2)^{(2)}$ flavor symmetries act respectively. The theory on the right-hand side is obtained by gauging the $SU(2)^{(1)}$ symmetry with $\mathcal{N}=3$ Chern-Simons terms with level $k=3$. The $\mathcal{N}=3$ theory is expected to flow to an $\mathcal{N}=4$ SCFT with a SUSY enhancement, due to the nilpotent property of the moment maps of the $SU(2)^{(1)}$ Gaiotto:2008ak . The enhancement can be explicitly confirmed using the superconformal index Evtikhiev:2017heo . Gauging lifts all the Coulomb branch, leaving the theory with only the Higgs branch $\mathbb{C}^{2}/\mathbb{Z}_{2}$ and $SU(2)$ flavor symmetry, which matches that of the $(A_{1},A_{3})$ EFT. This duality can also be verified through computations of the superconformal index, in particular one can detect the presence of the topological sector by turning on background magnetic fluxes for the $SU(2)$ flavor symmetry Beratto:2021xmn . The round 3-sphere partition functions $Z_{S^{3}}$ for the theories appearing in the duality can be computed using localization, yielding